OpenStudy (anonymous):

y = −4x + 2 What is the slope and Y intercept?

OpenStudy (anonymous):

slope= -4x Y-intercept= 2

OpenStudy (aravindg):

y=mx+c where m is slope and c is y intercept!!

OpenStudy (anonymous):

Use the slope and the y-intercept to graph y = −4x + 2 .... 

OpenStudy (anonymous):

What does the graphed line look like though? Im confused

OpenStudy (aravindg):

well what i do fo graphing is make it in form \[\large \frac{x}{a}+\frac{y}{b}=`1\]

OpenStudy (aravindg):

where a is x i ntercept b is y intercept

OpenStudy (aravindg):

the line joining these 3 intercepts will give you the line you need !!

OpenStudy (aravindg):

*2 intercepts